# Crypto Inverse-Power Options and Fractional Stochastic Volatility

### v2

@@ 5 Empirical analysis
In this section, we present a thorough demonstration of the efficacy of our FSV model through meticulous calibration exercises with Bitcoin options. Via comparative analyses against several benchmark models, we not only underscore the superior capacity and flexibility of our FSV model in effectively capturing the key traits of crypto market dynamics but unveil the substantial computational advantage of the type-III fractional kernel relative to other kernel types as well. Notably, to demonstrate the adeptness of the FSV model in accommodating the erratic dynamics of the highly speculative and unregulated crypto market, it is necessary to evaluate the model performance across different time periods. Towards that end, we draw upon two independent data sets: one representing abnormal market conditions amidst the COVID-19 pandemic, and the other reflecting current market dynamics. After the calibration, we then extract the calibrated parameters from the two best-performing models to further investigate the impact of the power mechanism on corresponding Quanto inverse options.
@@ 5.1 Data preparation
We collect tick-level trade data on Bitcoin options from the Debirit platform (deribit.com) to conduct the calibration exercises. All listed options are of European style. The first data set (labeled 1), directly available from [Xia, 2021, Sect. 4.2] [58], contains 40 call options traded as of July 11, 2020, covering four maturities $T=19,47,166,257$ days. Each maturity covers 10 options with different strike prices ($K$) ranging from \$5,000 to \$32,000. The second data set (labeled 2) is new, consisting of call options traded as of February 19, 2024, also under four maturities, $T=4,39,130,312$ days, with strike price range \$$\$[50,000,200,000]\ni K$. The Bitcoin spot prices ($S_{0}$) are also directly sourced from the Deribit platform and stand at \$9,232.98 and \$52,108.00 for the first and second data sets, respectively. Market prices are determined based on the last prices of the call options.
@@
    $C_{0}(K,T)\geq C_{0}(K^{\prime},T)$ if $K\leq K^{\prime}$, with $C_{0}(K,T)$ denoting the observe call option price with strike price $K$ and maturity $T$;
@@
This procedure only eliminates approximately 10% of price data, with the majority satisfying the above conditions. As a result, data set 2 also contains 40 call options, with each maturity including exactly 10 prices as in data set 1.
@@ 5.3 Calibration results and discussions
The calibration results are consolidated in Table 1, Table 2, Table 3, Table 4, Table 5, and Table 6, presenting the calibrated parameter values for the six implemented models individually for each data set, along with the minimized ARPEs. Wall time measurements for the FSV models using different kernel types are also included. [^18] All implementation programs are coded in MATLAB® and executed on a personal computer with a single 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz processor. All numerical results are rounded to six significant digits and the best results are marked as “ $\star$ .” Figure 3 and Figure 4 further provide visual comparisons between the calibrated model prices and the market prices, for which we only include the best-fitted FSV models and the benchmark models for clarity.
@@
On a first look at the calibration results, it is seen that the FSV models easily outperform the popular Black–Sholes model and Heston model as benchmarks – a direct evidence that correlated jumps in prices and volatility are essential for describing the crypto market dynamics. The observable levels of variability in the calibrated parameters under different types of fractional kernels are mostly due to these kernels exhibiting different local behaviors, despite the same asymptotic behaviors. In terms of model performance, regardless of the base processes in use, all three kernels have demonstrated comparable results, though the type III tends to generate lower ARPEs. Particularly striking is the significantly reduced wall time associated with the type-III kernel, which underscores its suitability for large-scale computational tasks. We stress that this enhanced efficiency comes from the type-III kernel’s ability to reexpress the numerical integral for $\log\phi_{Y_{1}}$ in (1) into a closed-form formula, which can be efficiently implemented thanks to Proposition 3. Moreover, the comparison of the FSV models against the benchmarks with usual stochastic volatility reveals that the former have consistently outperformed. This observation undoubtedly corroborates the presence of short-range dependence of stochastic volatility in the crypto market, as previously evidenced in [Takaishi, 2020] [49], underlying its practical necessity.
@@
On a closer look at the calibrated parameter values, we first see that the fraction parameter $\hat{d}$, directly linked to the degree of short-term dependency, is significantly different from 1 irrespective of the data set under consideration. As discussed in [Wang and Xia, 2022] [53], a higher value of $d$ closer to $1$ indicates “smoother” sample paths of volatility, while a lower $d$ closer to 0.5 signifies “rougher” paths. Hence, this finding suggests the prevalence of short-range volatility dependence for Bitcoin, and by extension, the broader crypto market, which is largely consistent with the empirical evidence presented in [Takaishi, 2020] [49]. Consequently, an important implication is that the proposed FSV framework alone possesses the capability to capture this rough-volatility characteristic in the dynamics of crypto prices. As for the mean reversion-speed parameter $\hat{\kappa}$, we observe from Table 1 and Table 2 that the models calibrated on data set 1 demonstrate moderate values ranging from 1 to 5, though the FSV-ALJD model under the type-III kernel comes with a higher value, around 9; in contrast, Table 4 and Table 5 generally reveal higher values of $\hat{\kappa}$ calibrated from data set 2 in comparison to the results obtained from the first. This disparity can be attributed to the increasing efficiency of the Bitcoin market in 2024, compared with the conditions prevailing during the COVID-19 pandemic in 2020, for a higher mean-reverting speed typically suggests a rapid dissemination of information and sentiment within the market. This outcome highlights the adeptness of FSV models in capturing the mean-reverting nature of Bitcoin volatility, a stylized fact across equity and commodity markets as well. Regarding the leverage parameter $\hat{\rho}$, which governs the jump correlation between crypto prices and volatility, our findings reveal a consistently positive moderate association across all stochastic-volatility models with jumps, which affirms the existence of the so-called “inverse leverage effect,” wherein a sudden increase in price coincides with a surge in volatility. It is crucial to note that this inverse leverage effect is particularly pronounced within jump components, when jump models are being used. On the other hand, the calibrated Heston model, which is entirely diffusive, still suggests negative price–volatility correlations ($\hat{\rho}$), indicating a direct leverage effect. [^19] An additional insight reveals that within the Heston model, the volatility-of-volatility scale ($\hat{\varsigma}$) can soar to as high as 1,500%, implying that Bitcoin volatility itself exhibits significant volatility, a notion consistent with the findings presented in [Madan et al., 2019, Sect. 3.4] [43]. This interesting observation aligns precisely with the empirical findings reported by [Huang et al., 2022] [29] using crypto time series data; see also [Hou et al, 2020, Tab. 1] [28]. It is also worth mentioning that this phenomenon is known to prevail in commodity markets ([Trolle and Schwartz, 2009] [51]) as well as other highly speculative markets such as the Chinese stock market ([Hou, 2013] [27]). Hence, a pertinent implication is the continued persistence of the inverse leverage effect within the Bitcoin market.
@@
Now we comment on the calibrated parameters specific to the base processes. In the ALJD case, particular attention is drawn to the diffusion volatility ($\hat{\sigma}_{X}$), the jump intensities ($\hat{\lambda}_{X}$ and $\hat{\lambda}_{Y}$), and the jump asymmetry parameter ($\hat{\eta}$). As can be seen from Table 1 and Table 4, The calibration results for both data sets indicate a significant volatility of asset returns, exceeding 50% on the annual basis. Furthermore, it is noted that the jump intensity of volatility (or instantaneous activity) surpasses that of the crypto prices, which suggests the significance of upward jumps in the volatility dynamics of cryptos. Regarding $\hat{\eta}$, it is observed that all optimized values deviate from 1, implying asymmetry in the distribution of Bitcoin price jumps. Specifically, nearly all implemented models yield $\hat{\eta}>1$, which indicates that downward return jumps are more aggressive than upward return jumps on average, except for the model utilizing the type-II kernel calibrated on data set 2. In the GMRTS case, note that with $c_{X}=1/2$, $X$ has sample paths of infinite variation, same as the one in the ALJD case where a diffusion component is present, while the parameter $n=2$ has no effect on the path regularity of $X$; we refer to [Fei and Xia, 2024, Sect. 6] [22] for a detailed discussion. In this instance, the negative value of the skewness parameter $\hat{\theta}$ validates the left-skewed nature of Bitcoin returns on average, which is precisely in line with the implications of the ALJD models. Also, the roles of the shape parameters $a_{X}$ and $a_{Y}$ mirror those of the intensity parameters in the ALJD scenario – they govern the levels of jump activity in the base process $X$ and volatility jump process $Y$ respectively, carrying practical implications akin to trading activity. From the tables, it is evident that $\hat{a}_{X}$ generally exceeds 1 significantly. When compared to $\hat{a}_{Y}$, this suggests that Bitcoin tends to experience heightened trading activity despite its elevated volatility and tail risks, which is consistent with the empirical findings presented in [Fei and Xia, 2024] [22] using time series data. Moreover, the rate parameter $\hat{b}_{X}$ simply acts as a multiplier, dictating the scale of price jumps.
@@ 5.4 Sensitivity analysis for Quanto inverse-power options
Using the optimized parameter values from the two models associated with the smallest ARPEs for the two data sets, identified as the FSV-GMRTS model under the type-III kernel and the FSV-ALJD model under the type-III kernel, respectively, we now explore the influence of imposing powers on crypto inverse options, in order to draw insights into the nonlinear leverage effect resulting from the utilization of the power mechanism. Simultaneously, we aim to showcase the general characteristic function-based pricing formulas outlined in Proposition 2, thereby extending the pricing methodology initially proposed by [Alexander et al., 2023] [2] to encompass general stochastic-volatility models. For specificity, in each dataset, we set the conversion rate equal to the spot Bitcoin/USD rate, $R=S_{0}$, and focus on the nearest strike price to $S_{0}$ and the second-longest maturity. The near-the-money Quanto inverse-power option prices are then determined by the model prices obtained from the two best-performing models.
@@ 6 Conclusions
Since the majority of actively traded crypto options are of the inverse type through the Deribit platform, USD-denominated investors naturally face significant currency risk when holding these options. This risk can be effectively mitigated through the utilization of Quanto inverse options, as introduced by [Alexander et al., 2023] [2], by employing predetermined conversion rates. While emphasizing the essential functionalities of Quanto inverse options, the present paper also introduces generalizations featuring nonlinear power payoffs. The resulting Quanto inverse-power options offer investors the flexibility to further adjust their exposure to the inherent risk of cryptos. It is important to note that all Quanto inverse-power options are valued in fiat currencies like USD. We have developed (in Section 3) efficient Fourier transform-type pricing–hedging formulas for this broader class of exotic crypto options, all formulas being model-independent and requiring none but a straightforward characteristic function. Noteworthily, these formulas are presented in a conditional format, allowing for rapid implementation in dynamic pricing and hedging scenarios where frequent model updating may be undesirable.
@@
Appendix Appendix A Mathematical proofs
@@
$=\frac{S^{{\rm i}u}_{t_{0}}}{\phi^{{\rm i}u(t-t_{0})}_{Y_{1}}(-{\rm i}\rho)}\mathsf{E}_{t_{0}}\big{[}\mathsf{E}\big{[}e^{{\rm i}u(X_{\mathcal{T}_{t}}-X_{\mathcal{T}_{t_{0}}}+\rho(Y_{t}-Y_{t_{0}})-\log\phi_{X_{1}}(-{\rm i})(\mathcal{T}_{t}-\mathcal{T}_{t_{0}}))}\big{|}\upsigma(Y_{s}:s\in[0,t])\vee\mathscr{F}_{t_{0}}\big{]}\big{]}$
@@
where the last equality follows from the infinite divisibility of $\mathcal{L}(X_{1})$. We remark that $\mathbb{F}$ contains the $\mathcal{T}$-stopped filtration of $X$, which is well-defined because of the path continuity of $\mathcal{T}$. By substituting the expression for $\mathcal{T}$ in terms of $Y$ and writing $B(t_{0},t)=(A_{0}-m)(e^{-\kappa t_{0}}-e^{-\kappa t})/\kappa+m(t-t_{0})$, the exponent in the last equality in (Appendix A) is equal to
@@
where the first equality uses (Appendix A). The last result is the sum of two terms, one independent of $\mathscr{F}_{t_{0}}$ and the other measurable with respect to $\mathscr{F}_{t_{0}}$. Then, using further the infinite divisibility of $\mathcal{L}(Y_{1})$ we have for $u\in\mathbb{R}$,
@@
$\mathsf{E}_{t_{0}}[[X_{\mathcal{T}},X_{\mathcal{T}}]_{t}-[X_{\mathcal{T}},X_{\mathcal{T}}]_{t_{0}}]$ $=\mathsf{E}_{t_{0}}[\mathsf{E}_{t_{0}}[[X,X]_{\mathcal{T}_{t}}-[X,X]_{\mathcal{T}_{t_{0}}}|\upsigma(Y_{s}:s\in[0,t])\vee\mathscr{F}_{t_{0}}]]$
@@
For the cross-variation, the claim is that $[X_{\mathcal{T}},Y]_{t}=0$, $\mathsf{Q}$-a.s., despite the dependence between $\mathcal{T}$ and $Y$. To see this, recall that $\mathcal{T}$ is a strictly increasing process by (2.1.1) and the strict positivity of $A$, so that $\mathcal{T}^{-1}$ exists $\mathsf{Q}$-a.s. and is also strictly increasing, thus being a valid time change. Reusing the previous argument based on path continuity, we have that $[X_{\mathcal{T}},Y]_{t}=[X,Y_{\mathcal{T}^{-1}}]_{\mathcal{T}_{t}}$ $\mathsf{Q}$-a.s., but since $t\in(0,T]$ is arbitrary and $[X,Y_{\mathcal{T}^{-1}}]$ has càdlàg sample paths by construction, this means that $[X_{\mathcal{T}},Y]_{t}=0$ if and only if $[X,Y_{\mathcal{T}^{-1}}]_{t}=0$. As $Y$ is purely discontinuous, it suffices to consider the case where $X$ is also purely discontinuous, or has no Brownian part, with
@@
Lastly, note that $\mathsf{E}[[X,X]_{1}]=\mathsf{Var}[X_{1}]$ and $\mathsf{E}[[Y,Y]_{1}]=\mathsf{Var}[Y_{1}]$ because $X$ and $Y$ are both Lévy processes. Putting (Appendix A) back in (Appendix A) gives the desired result upon rearrangement. ∎
@@
Based on (2.2.1), the distribution $\mathcal{L}(\log S_{t})$ is also absolutely continuous by the continuous mapping theorem. Using the conditional representation in (Appendix A) directly yields the continuity property for the conditional distribution on $\mathscr{F}_{t_{0}}$, given $t_{0}\in[0,t)$. ∎
@@
As before, fix $t_{0}\in[0,T)$ and let $g_{\log S_{T}|t_{0}}(x)$, $x\in\mathbb{R}$, denote the cumulative distribution function of $\log S_{T}$ conditional on $\mathscr{F}_{t_{0}}$, which also has a density function $f_{\log S_{T}|t_{0}}(x)$, $x\in\mathbb{R}$, due to Theorem 1.
@@
By treating the value of inverse power call option as a function of time $t$, the contemporaneous cryptocurrency price $S$, and the variance swap (inception) price $V_{S}\equiv V_{S}(\imath,T)$, namely $C^{\rm(qip)}_{t}\equiv C^{\rm(qip)}(t,S,V)$, we have by using Itô’s formula again that
@@
for $t\in[0,T]$. According to the pricing formulas in Proposition 1, assuming the validity of doing so, taking the corresponding partial derivatives of (3.1.9) comes down to differentiating the conditional characteristic function in (1) and in (Appendix A). After some simplifications we obtain that for every fixed $u\in\mathbb{R}$,

### v3

@@ 5 Empirical analysis
In this section, we present a thorough demonstration of the efficacy of our FSV model through meticulous calibration exercises with Bitcoin options. Via comparative analyses against several benchmark models, we not only underscore the superior capacity and flexibility of our FSV model in effectively capturing the key traits of crypto market dynamics but unveil the substantial computational advantage of the type-III fractional kernel relative to other kernel types as well. Notably, to demonstrate the adeptness of the FSV model in accommodating the erratic dynamics of the highly speculative crypto market, it is necessary to evaluate the model performance across different time periods. Towards that end, we draw upon two independent data sets: one representing abnormal market conditions amidst the COVID-19 pandemic, and the other reflecting recent market dynamics. After the calibration, we then extract the calibrated parameters from the two best-performing models to further investigate the impact of the power mechanism on corresponding Quanto inverse options.
@@ 5.1 Data preparation
We collect tick-level trade data on Bitcoin options from the Debirit exchange (deribit.com) to conduct the calibration exercises. All listed options are of European style. The first data set (labeled 1), directly available from [Xia, 2021, Sect. 4.2] [57], contains 40 call options traded as of July 11, 2020, covering four maturities $T=19,47,166,257$ days. Each maturity covers 10 options with different strike prices ($K$) ranging from \$5,000 to \$32,000. The second data set (labeled 2) is new, consisting of call options traded as of February 19, 2024, also under four maturities, $T=4,39,130,312$ days, with strike price range \$$\$[50,000,200,000]\ni K$. The Bitcoin spot prices ($S_{0}$) are also directly sourced from the Deribit exchange and stand at \$9,232.98 and \$52,108.00 for the first and second data sets, respectively. Market prices are determined based on the last prices of the calls.
@@
    $C_{0}(K,T)\geq C_{0}(K^{\prime},T)$ if $K\leq K^{\prime}$, with $C_{0}(K,T)$ denoting the observed call price with strike price $K$ and maturity $T$;
@@
This procedure only eliminates approximately 10% of price data, with the majority satisfying the above conditions. As a result, data set 2 also contains 40 calls, with each maturity including exactly 10 prices as in data set 1.
The two data sets were intentionally selected to reflect structurally different market conditions. As can be seen in the forthcoming Figure 3 and Figure 4, data set 1 capturing the COVID-19 period exhibits a pronounced clustering of option prices across comparable maturities, which pattern is indicative of elevated volatility-of-volatility and substantially weaker mean reversion relative to data set 2 corresponding to the post-COVID period. Comprehensibly, the compression of the term structure can be attributed to the significant elevation of short-term uncertainty such that longer maturities offered little additional protection, indicating the extreme levels of market instability at that time in the crypto market.
@@ 5.3 Calibration results and discussions
The calibration results are consolidated in Table 1, Table 2, Table 3, Table 4, Table 5, and Table 6, presenting the calibrated parameter values for the six implemented models individually for each data set, along with the minimized ARPEs. Wall time measurements for the FSV models using different kernel types are also included. [^19] All calibration programs are coded in MATLAB® and executed on a personal computer with a single 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz processor. All numerical results are rounded to six significant digits and the best results are marked as “ $\star$ .” Figure 3 and Figure 4 provide visual comparisons between the calibrated model prices and the market prices, for which we only include the best-fitted FSV models and the benchmark models for clarity. Figure 5 further compares the individual relative pricing errors (RPEs) (namely the summands in (5.2.1)) across the best-fitted FSV models, the Heston model, and the Black–Scholes model.
@@
Data set 1 (July 11, 2020)
Data set 2 (February 19, 2024)
Figure 5: Comparison of relative pricing errors
@@
On a first look at the calibration results, it is seen that the FSV models easily outperform the popular Black–Sholes model and Heston model as benchmarks – a direct evidence that correlated jumps in prices and volatility are essential for describing the crypto market dynamics. The observable levels of variability in the calibrated parameters under different types of fractional kernels are mostly due to these kernels exhibiting different local behaviors, despite the same asymptotic behaviors. In terms of model performance, regardless of the base processes in use, all three kernels have demonstrated comparable results, though the type III tends to generate lower ARPEs. Particularly striking is the significantly reduced wall time associated with the type-III kernel, which underscores its suitability for large-scale computational tasks. We stress that this enhanced efficiency comes from the type-III kernel’s ability to reexpress the numerical integral for $\log\phi_{Y_{1}}$ in (1) into a closed-form formula, which can be efficiently implemented thanks to Proposition 3. Moreover, the comparison of the FSV models against the benchmarks with usual stochastic volatility reveals that the former have consistently outperformed. This observation undoubtedly corroborates the presence of short-range dependence of stochastic volatility in the crypto market, as previously evidenced in [Takaishi, 2020] [48], underlying its practical necessity. Furthermore, as seen from Figure 5, for the two data sets, respectively, the maximum relative pricing errors of the FSV models stand at around 15% and 29%, significantly lower than those of the Black–Scholes model (100% and 96%) and the Heston model (77% and 86%). In particular, the FSV models have demonstrated superior performance on data with strike prices above \$10,000 and \$100,000 for the two data sets, respectively, indicating its strong capability in accurately pricing deeply out-of-the-money options.
@@
On a closer look at the calibrated parameters, we see that the fraction parameter $\hat{d}$, directly linked to the degree of short-term dependency, is significantly different from 1 irrespective of the data set under consideration. As discussed in [Wang and Xia, 2022] [52], a higher value of $d$ closer to $1$ indicates “smoother” sample paths of volatility, while a lower $d$ closer to 0.5 signifies “rougher” paths. Hence, this finding suggests the prevalence of short-range volatility dependence for Bitcoin, and by extension, the broader crypto market, which is largely consistent with the empirical evidence presented in [Takaishi, 2020] [48]. Consequently, an important implication is that the proposed FSV framework alone possesses the capability to capture this rough-volatility characteristic in the dynamics of crypto prices. As for the mean reversion-speed parameter $\hat{\kappa}$, we observe from Table 1 and Table 2 that the models calibrated on data set 1 demonstrate moderate values ranging from 1 to 5, though the FSV-ALJD model under the type-III kernel comes with a higher value, around 9; in contrast, Table 4 and Table 5 generally reveal higher values of $\hat{\kappa}$ calibrated from data set 2 in comparison to the results obtained from the first. This disparity can be attributed to the increasing efficiency of the Bitcoin market in 2024, compared with the conditions prevailing during the COVID-19 pandemic in 2020, for a higher mean-reverting speed typically suggests a rapid dissemination of information and sentiment within the market. This outcome highlights the adeptness of FSV models in capturing the mean-reverting nature of Bitcoin volatility, a stylized fact across equity and commodity markets as well. Regarding the leverage parameter $\hat{\rho}$, which governs the jump correlation between crypto prices and volatility, our findings reveal a consistently positive moderate association across all stochastic-volatility models with jumps, which affirms the existence of the so-called “inverse leverage effect,” wherein a sudden increase in price coincides with a surge in volatility. It is crucial to note that this inverse leverage effect is particularly pronounced within jump components, when jump models are being used. On the other hand, the calibrated Heston model, which is entirely diffusive, still suggests negative price–volatility correlations ($\hat{\rho}$), indicating a direct leverage effect. [^20] Another observation reveals that within the Heston model, the volatility-of-volatility scale ($\hat{\varsigma}$) can soar to as high as 1,500%, implying that Bitcoin volatility itself exhibits significant volatility, a notion consistent with the findings presented in [Madan et al., 2019, Sect. 3.4] [42]. This interesting observation aligns precisely with the empirical findings reported by [Huang et al., 2022] [29] using crypto time series data; see also [Hou et al, 2020, Tab. 1] [28]. It is also worth mentioning that this phenomenon is known to prevail in commodity markets ([Trolle and Schwartz, 2009] [50]) as well as other highly speculative markets such as the Chinese stock market ([Hou, 2013] [27]). Hence, a pertinent implication is the continued persistence of the inverse leverage effect within the Bitcoin market.
@@
Now we look at the calibrated parameters specific to the base processes. In the ALJD case, particular attention is drawn to the diffusion volatility ($\hat{\sigma}_{X}$), the jump intensities ($\hat{\lambda}_{X}$ and $\hat{\lambda}_{Y}$), and the jump asymmetry parameter ($\hat{\eta}$). As can be seen from Table 1 and Table 4, The calibration results for both data sets indicate a significant volatility of asset returns, exceeding 50% on the annual basis. Furthermore, it is noted that the jump intensity of volatility (or instantaneous activity) surpasses that of the crypto prices, which suggests the significance of upward jumps in the volatility dynamics of cryptos. Regarding $\hat{\eta}$, it is observed that all optimized values deviate from 1, implying asymmetry in the distribution of Bitcoin price jumps. Specifically, nearly all implemented models yield $\hat{\eta}>1$, which indicates that downward return jumps are more aggressive than upward return jumps on average, except for the model utilizing the type-II kernel calibrated on data set 2. We also observe a substantial difference between the estimates $\hat{\lambda}_{X}$ and $\hat{b}_{X}$, which correspond to jump intensity and (inverse) jump scale, respectively, across the two data sets. In particular, comparing the 2024 period to that of 2020, the estimate $\hat{b}_{X}$ has increased notably, which suggests a potential narrowing in the distribution of jump sizes; at the same time, the higher value of $\hat{\lambda}_{X}$ indicates an increased frequency of return jumps. In other words, the crypto market appears to be experiencing more frequent but smaller jumps, and this shift may signal a more stable trading environment – one marked by greater market participation and a predominance of milder informational shocks, in contrast to the heightened uncertainty and large jump magnitudes observed during the COVID-19 period.
In the GMRTS case, note that with $c_{X}=1/2$, $X$ has sample paths of infinite variation, same as the one in the ALJD case where a diffusion component is present, while the parameter $n=2$ has no effect on the path regularity of $X$; we refer to [Fei and Xia, 2024, Sect. 6] [22] for a detailed discussion. In this instance, the negative value of the skewness parameter $\hat{\theta}$ validates the left-skewed nature of Bitcoin returns on average, which is precisely in agreement with the implications of the ALJD models. Also, the roles of the shape parameters $a_{X}$ and $a_{Y}$ mirror those of the intensity parameters in the ALJD scenario – they govern the levels of jump activity in the base process $X$ and volatility jump process $Y$ respectively, carrying practical implications akin to trading activity. From the tables, it is evident that $\hat{a}_{X}$ generally exceeds 1 significantly. When compared to $\hat{a}_{Y}$, this suggests that Bitcoin tends to experience heightened trading activity despite its elevated volatility and tail risks, which is consistent with the empirical findings presented in [Fei and Xia, 2024] [22] using time series data. Moreover, the rate parameter $\hat{b}_{X}$ simply acts as a multiplier dictating the scale of price jumps.
To further illustrate the superior performance of the proposed FSV models in capturing the key dynamics, Figure 6 presents the implied volatility surfaces derived from the Bitcoin option price data and the corresponding best-fitted models (Table 2 and Table 4), identified as the FSV-GMRTS model under the type-I kernel and the FSV-ALJD model under the type-III kernel for the two data sets, respectively. The volatility surfaces are constructed via two-dimensional linear interpolation based on 40 implied volatilities across all strike prices and maturities in each data set. Note that the wide range of implied volatilities (exceeding 200% in some price regions) reflects the extreme volatility of Bitcoin over the observed periods and, more importantly, despite the markedly different shapes of the implied volatility surfaces (in line with dissimilar market conditions as noted earlier), the best-fitted models successfully replicate the observed surfaces with high accuracy – particularly in the out-of-the-money regions – underlying the FSV models robustness across different strike price and maturity levels.
Data set 1 (July 11, 2020)
Data set 2 (February 19, 2024)
Figure 6: Comparison of market and model implied volatility surfaces
@@ 5.4 Sensitivity analysis for Quanto inverse-power options
Using the optimized parameter values from the two models associated with the smallest ARPEs for the two data sets, the FSV-GMRTS model under the type-I kernel and the FSV-ALJD model under the type-III kernel, respectively, we now explore the influence of imposing powers on crypto inverse options, in order to draw insights into the nonlinear leverage effect resulting from the utilization of the power mechanism. Simultaneously, we aim to showcase the general characteristic function-based pricing formulas outlined in Proposition 2, thereby extending the pricing methodology initially proposed by [Alexander et al., 2023] [2] to encompass general stochastic-volatility models. For specificity, in each dataset, we set the conversion rate equal to the spot Bitcoin/USD rate, $R=S_{0}$, and focus on the nearest strike price to $S_{0}$ and the second-longest maturity. The near-the-money Quanto inverse-power option prices are then determined by the model prices obtained from the two best-performing models.
@@ 6 Conclusions
Since the majority of actively traded crypto options are of the inverse type through the Deribit exchange, USD-denominated investors naturally face significant currency risk when holding these options. This risk can be effectively mitigated through the utilization of Quanto inverse options, as introduced by [Alexander et al., 2023] [2], by employing predetermined conversion rates. While emphasizing the essential functionalities of Quanto inverse options, the present paper also introduces generalizations featuring nonlinear power payoffs. The resulting Quanto inverse-power options offer investors the flexibility to further adjust their exposure to the inherent risk of cryptos. It is important to note that all Quanto inverse-power options are valued in fiat currencies like USD. We have developed (in Section 3) efficient Fourier transform-type pricing–hedging formulas for this broader class of exotic crypto options, all formulas being model-independent and requiring none but a straightforward characteristic function. Noteworthily, these formulas are presented in a conditional format, allowing for rapid implementation in dynamic pricing and hedging scenarios where frequent model updating may be undesirable.
@@
Appendix A Mathematical proofs
@@
$=\frac{S^{{\rm i}u}_{t_{0}}}{\phi^{{\rm i}u(t-t_{0})}_{Y_{1}}(-{\rm i}\rho)}\mathsf{E}_{t_{0}}\big{[}\mathsf{E}\big{[}e^{{\rm i}u(X_{\mathcal{T}_{t}}-X_{\mathcal{T}_{t_{0}}}+\rho(Y_{t}-Y_{t_{0}})-\log\phi_{X_{1}}(-{\rm i})(\mathcal{T}_{t}-\mathcal{T}_{t_{0}}))}\big{|}\upsigma(Y_{s}:s\in[0,t])\vee\mathcal{F}_{t_{0}}\big{]}\big{]}$
@@
where the last equality follows from the infinite divisibility of $\mathcal{L}(X_{1})$. We remark that $\mathbb{F}$ contains the $\mathcal{T}$-stopped filtration of $X$, which is well-defined because of the path continuity of $\mathcal{T}$. By substituting the expression for $\mathcal{T}$ in terms of $Y$ and writing $B(t_{0},t)=(A_{0}-m)(e^{-\kappa t_{0}}-e^{-\kappa t})/\kappa+m(t-t_{0})$, the exponent in the last equality in (A) is equal to
@@
where the first equality uses (A). The last result is the sum of two terms, one independent of $\mathcal{F}_{t_{0}}$ and the other measurable with respect to $\mathcal{F}_{t_{0}}$. Then, using further the infinite divisibility of $\mathcal{L}(Y_{1})$ we have for $u\in\mathbb{R}$,
@@
$\mathsf{E}_{t_{0}}[[X_{\mathcal{T}},X_{\mathcal{T}}]_{t}-[X_{\mathcal{T}},X_{\mathcal{T}}]_{t_{0}}]$ $=\mathsf{E}_{t_{0}}[\mathsf{E}_{t_{0}}[[X,X]_{\mathcal{T}_{t}}-[X,X]_{\mathcal{T}_{t_{0}}}|\upsigma(Y_{s}:s\in[0,t])\vee\mathcal{F}_{t_{0}}]]$
@@
For the cross-variation, the claim is that $[X_{\mathcal{T}},Y]_{t}=0$, $\mathsf{Q}$-a.s., despite the dependence between $\mathcal{T}$ and $Y$. To see this, recall that $\mathcal{T}$ is a continuous strictly increasing process by (2.1.1) and the strict positivity of $A$, so that $\mathcal{T}^{-1}$ exists $\mathsf{Q}$-a.s. and is also strictly increasing, thus being a valid time change. Reusing the previous argument based on path continuity, we have that $[X_{\mathcal{T}},Y]_{t}=[X,Y_{\mathcal{T}^{-1}}]_{\mathcal{T}_{t}}$ $\mathsf{Q}$-a.s., but since $t\in(0,T]$ is arbitrary and $[X,Y_{\mathcal{T}^{-1}}]$ has càdlàg sample paths by construction, this means that $[X_{\mathcal{T}},Y]_{t}=0$ if and only if $[X,Y_{\mathcal{T}^{-1}}]_{t}=0$. As $Y$ is purely discontinuous, it suffices to consider the case where $X$ is also purely discontinuous, or has no Brownian part, with
@@
Lastly, note that $\mathsf{E}[[X,X]_{1}]=\mathsf{Var}[X_{1}]$ and $\mathsf{E}[[Y,Y]_{1}]=\mathsf{Var}[Y_{1}]$ because $X$ and $Y$ are both Lévy processes. Putting (A) back in (A) gives the desired result upon rearrangement. ∎
@@
Based on (2.2.1), the distribution $\mathcal{L}(\log S_{t})$ is also absolutely continuous by the continuous mapping theorem. Using the conditional representation in (A) directly yields the continuity property for the conditional distribution on $\mathcal{F}_{t_{0}}$, given $t_{0}\in[0,t)$. ∎
@@
As before, fix $t_{0}\in[0,T)$ and let $g_{\log S_{T}|t_{0}}(x)$, $x\in\mathbb{R}$, denote the cumulative distribution function of $\log S_{T}$ conditional on $\mathcal{F}_{t_{0}}$, which also has a density function $f_{\log S_{T}|t_{0}}(x)$, $x\in\mathbb{R}$, due to Theorem 1.
@@
By treating the value of inverse power call option as a function of time $t$, the contemporaneous cryptocurrency price $S$, and the variance swap (inception) price $V_{S}\equiv V_{S}(\imath,T)$, namely $C^{\rm(qip)}_{t}\equiv C^{\rm(qip)}(t,S,V_{S})$, we have by using Itô’s formula again that
@@
for $t\in[0,T]$. According to the pricing formulas in Proposition 1, assuming the validity of doing so, taking the corresponding partial derivatives of (3.1.9) comes down to differentiating the conditional characteristic function in (1) and in (A). After some simplifications we obtain that for every fixed $u\in\mathbb{R}$,